OpenStudy (anonymous):

Which of the following is the slope between the points (0, -10) and (-4, 0)? a. 2/5 b.-2/5 c.-5/2 d.5/2

OpenStudy (anonymous):

Do you know how to find the slope between 2 points?

hartnn (hartnn):

know the formula to find slope ?

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

hartnn (hartnn):

so what slope did u get ?

OpenStudy (anonymous):

I don't understand slope can you explain how i am supposed to set up the equation and then i could solve it

hartnn (hartnn):

First point , (0,-10) compare it with (x1,y1) u get x1=0, y1=-10 in same way find x2 and y2 and put all these 4 values in the formula i gave you to find slope =m

OpenStudy (anonymous):

so would it be 0,-4/-10,0 for the equation?

hartnn (hartnn):

the numerator has y-values

hartnn (hartnn):

so in the numerator, its -10-0 got this ? what will be in the denominator then ?

OpenStudy (anonymous):

the denominator would be -4-0 ?

OpenStudy (anonymous):

i meant 0-4 ?

hartnn (hartnn):

u missed a - *minus* sign \(\huge 0-(-4)= 0+4=? \)

hartnn (hartnn):

one minus sign comes from formula and one from the fact that the co-ordinate itself is negative , its -4

OpenStudy (anonymous):

so would the equation be -10-0/0+4

hartnn (hartnn):

yes! correct :) m=(-10-0)/(0+4)

hartnn (hartnn):

so m= ?

OpenStudy (anonymous):

hold on im working out the problem now

OpenStudy (anonymous):

so the answer would be -5/2

hartnn (hartnn):

yes, that is absolutely correct! good :)

OpenStudy (anonymous):

ok thank you so much :)

hartnn (hartnn):

welcome ^_^